Below are the processing costs of Well Inc. observed for producing a type of vitamin C tablets in bottles over the last 6 months:
Month
Units of bottles Produced
Processing Cost
1
9500
$34,000
2
13,000
$42,000
3
17,000
$49,000
4
7,500
$28,000
5
11,000
$37,000
6
8,500
$30,000
a. Use the high low method to determine the variable cost per unit and fixed cost per month. Express your results in the form of a linear equation
y = a + bx. ​​​​​​​​(15 marks)
b. Predict the total processing cost if 22,000 units are produced. (10 marks)
Answer #1

a.

Variable cost per unit = (Highest activity cost – Lowest activity cost)/(Highest activity unit – Lowest activity unit)

= (49,000 - 28,000)/(17,000 - 7,500)

= 21,000/9,500

= $2.21

Fixed cost = Highest activity cost – (Variable cost per unit x Highest activity unit)

= 49,000 - (2.21 x 17,000)

= 49,000 - 37,570

= $11,430

b.

Total cost when 22,000 units are produced or used = Fixed cost + (Variable cost per unit x Number of units produced)

= 11,430 + (2.21 x 22,000)

= 11,430 + 48,620

= $60,050